Yesterday and today I was in the lab doing many cavity scan.
First I did many measurement with the cavity aligned in order to get the position of the 00 modes, then I misaligned the beam in many different ways to enhance the higher order modes.
In particular, I first misaligned the mode cleaner to make the beam clipping into the Faraday. To do this, I set to 0 the WFS gain, but I left the autolocker still enabled. In this way, the autolocker couldn't bring the mirrors back to the aligned position.
Then I misaligned also the TT2 to get even more HOMs.
Eventually, Rana came and we misaligned TT1 to clip the beam, and using TT2 we aligned back the beam to the arm.
To increase the SNR, we changed the gain of the TRY PD, setting it to 20dB (which corresponds to a factor 100 in digital scale)
I attached one scan that I did with Rana on Sunday night. I could not upload a better resolution image because the file size was too big, but here's the path to find all of the scans:
There are many folders, one per each day I measured. In each folder there are measurements relative to aligned cavity, Pitch and Yaw misalignment.
The PDA520 used for TRY was set to 0 dB analog gain. This corresponds to ~500 counts out of 32768. The change to 20 dB actually increases the gain by 100. This makes the single arm lock saturate at ~25000 counts (obviously in analog before the ADC). The right setting for our usual running is probably 10 dB.
For the IMC WFS, we had disabled the turn on in the autolocker to use the IMC to steer the beam in the FI, but that was a flop (not enough range, not enough lever arm). In the end, I think we didn't get any clipping.